# Visual Planning: Let’s Think Only with Images

### v2

@@ 1 Introduction
Large Language Models (LLMs) [7, 41, 2] have demonstrated strong capabilities in language understanding and generation, as well as growing competence in complex reasoning, enabled by their chain-of-thought reasoning abilities [54]. Building on these advances, recent work extends LLMs to support multiple modalities, yielding so-called Multimodal Large Language Models (MLLMs) [46, 25]: they incorporate visual embedded information at the input to tackle a broader spectrum of tasks, such as visual spatial reasoning [36, 33] and navigation [16, 32]. However, despite their multimodal inputs, these methods perform reasoning purely in the text format during inference, from captioning visual content [22] to generating verbal rationales [63].
@@
Building on this observation, we argue that performing multimodal reasoning only in the text pathway may not always offer the most intuitive or effective strategy, particularly for tasks that depend heavily on visual information and/or are ‘vision-first’ by design. Indeed, recent results from multimodal benchmarks [47, 33, 8, 9] offer growing evidence that purely language-based reasoning falls short in certain domains, particularly those involving spatial, geometric, or physical dynamics [60]. Such reliance on grounding visual information into text before reasoning introduces a modality gap that hinders the model’s ability to capture visual features and state transitions. This highlights a potential shortcoming of current MLLMs: while they process image inputs, they do not naturally “think” in images. For instance, tasks such as planning a route through a maze, designing the layout of a room, or predicting the next state of a mechanical system are often better served by visual representations, as verbal descriptions may be less effective and struggle to accurately capture complex spatial reasoning relationships. These examples suggest a broader question, which we aim to tackle in this work: can models directly plan in non-verbal modalities, such as images, without being mediated by text?
@@
Cognitive science also offers compelling motivation for this question [39]. Dual Coding Theory [42] proposes that human cognition operates through both verbal and nonverbal channels, each capable of independent representational and inferential processes. Recent work on MLLMs incorporates interleaved text and images as reasoning steps [24, 34]. However, they still remain fundamentally text-driven and rely on tool-based visualizations as auxiliary information for reasoning traces, with reasoning still mainly embedded in verbal traces. For instance, Visual Sketchpad [24] employs external tools to generate sketches as visual aids, and MVoT [34] generates per-step visualizations from language-based actions but still reasons in text for decision-making. As such, a truly visual-only reasoning paradigm that avoids any text-based reasoning proxies remains underexplored.
@@
In this work, we propose a new paradigm, Visual Planning, where reasoning is structured as a sequence of images, but without the mediation of language. To the best of our knowledge, this is the first attempt to investigate whether models can achieve planning purely through visual representations. Rather than generating textual rationales and answers, our approach produces step-by-step visualizations that encode planning or inference steps directly in images. As a pioneering exploration, it circumvents the modality mismatch that occurs when visual problems must be forced into explanations in verbal form, reinforces state transitions, and provides a new trackable interface for tasks like navigation [33], and visual problem-solving [22].
@@
Specifically, we explore this paradigm using the Large Vision Model (LVM) [4] trained exclusively on images and video frames with no textual data. This design choice removes potential confounders introduced by language-based supervision and enables a clean investigation of whether models can reason purely within the visual modality. Motivated by the success of reinforcement learning in acquiring reasoning capabilities within the language modality [17] and its strong generalization performance [12], we propose Visual Planning via Reinforcement Learning (VPRL), a novel two-stage reinforcement learning framework empowered by GRPO [48] for visual planning. It involves a distinct initializing stage for encouraging the exploration of the policy model in the given environment, which is then followed by reinforcement learning with a progress reward function.
@@
We validate the feasibility of our paradigms on grid-based navigation as a representative of spatial planning tasks, including Maze [26], FrozenLake [57], and MiniBehavior [28], where one agent is requested to navigate to a target location successfully without violating environment constraints. Our experiments reveal that the visual planning paradigm substantially surpasses the traditional textual reasoning method by supervised fine-tuning (SFT), achieving 27% higher average exact-match rate. In addition to better performance, our novel method VPRL exhibits stronger generalization to out-of-distribution scenarios than the SFT method in the visual planning paradigm (VPFT). To the best of our knowledge, we are the first to apply RL to image generation in the context of planning, with contributions as follows:
@@ 2.1 The Visual Planning Paradigm
The majority of prior visual reasoning benchmarks [15, 1, 59] can be and is typically tackled by grounding the visual information in the textual domain [19, 43, 61], followed by a few steps of textual reasoning. However, once the visual content is mapped to text (e.g., object names, attributes, or relations), the problem gets reduced to a language reasoning task, where the reasoning is carried out by the language model, even without reflecting any information from the visual modality.
@@ 2.2 Reinforcement Learning for Large Vision Models
Reinforcement learning (RL) has shown notable advantages in improving the generalization of autoregressive models by optimizing with sequence-level rewards beyond token-level supervision signals [12]. In autoregressive image generation, an image is represented as a sequence of visual tokens. Inspired by the success of RL in language reasoning [17], we introduce an RL-based training framework for visual planning empowered by Group Relative Policy Optimization (GRPO) [48]. It leverages the transitions between visual states to compute the reward signals while verifying the constraints from the environments. To enforce the policy model that generates valid actions with diverse exploration during the RL process, we then propose a novel two-stage reinforcement learning framework for visual planning. In Stage 1, we first apply supervised learning to initialize the policy model with random trajectories. Model’s visual planning is then optimized by the RL training in Stage 2.
@@
The dynamics interpreter $\mathcal{D}$ evaluates the transitions $a\in\mathcal{A}\cup\mathcal{E}$ for validity, which, by implementation, can be a dynamics model [44] or a rule-based system to elicit actions from state pairs, or a neural model as holistic validator that judges transitions without explicitly inferring actions. The progress estimator $P(v)$ quantifies progress by estimating the remaining steps or effort required to reach the goal from each visual state. By comparing the agent’s current and predicted state, we partition the generated candidate states $\mathcal{A}\cup\mathcal{E}$ into three disjoint subsets:
@@ 3 Experiments and Results
Tasks. To evaluate our proposed visual planning paradigm, we select representative tasks where planning can be expressed and executed entirely in the visual modality. We focus on tasks where state transitions are visually observable, distinguishing them from language-centric tasks like code generation [30] or traditional visual question answering. This design allows us to analyze planning behavior without relying on textual rationales or symbolic outputs. To compare visual planning with language-based reasoning, we experiment with 3 visual navigation environments: FrozenLake [57], Maze [26], and MiniBehavior [28]. All of them can be solved in both modalities, which enables a direct parallel comparison of pros and cons between visual planning and language reasoning strategies.
  • FrozenLake: It is initially proposed by Wu et al. [57] and implemented with Gym [6]. It simulates a grid-based frozen lake, where the agent is supposed to start from the designated position and find its way to the destination safely without falling into the ‘holes’.
@@
We construct synthetic datasets for the tasks with varying levels of complexity in patterns and environments. Details on data collection and implementation are provided in Appendix B.1.
Table 1: Performance of the closed- and open-source models on FrozenLake, Maze, and MiniBehavior. VPRL performs consistently the best (bold) across all tasks. denotes the post-trained model. ˜ represents texts and  represents images. The last column Avg. reports the average performance across three tasks.
Model Input Output FrozenLake Maze MiniBehavior Avg. EM (%) PR (%) EM (%) PR (%) EM (%) PR (%) EM (%) PR (%) Closed-Source Model Gemini 2.0 Flash   xxx- Direct ˜+ ˜ 21.2 47.6 8.3 31.4 0.7 29.8 10.1 36.3   xxx- CoT ˜+ ˜ 27.6 52.5 6.9 29.8 4.0 31.2 12.8 37.8 Gemini 2.5 Pro (think) ˜+ ˜ 72.0 85.0 21.5 35.5 37.6 59.9 43.7 60.1 Open-Source Model Qwen 2.5-VL-Instruct-7B   xxx- Direct ˜+ ˜ 1.2 15.0 0.6 14.5 0.3 9.8 0.7 13.1   xxx- CoT ˜+ ˜ 8.2 29.1 2.3 15.2 0.5 14.7 3.7 19.7   xxx- SFT ˜+ ˜ 68.6 84.4 60.9 70.3 31.3 56.1 53.6 69.9 LVM-7B   xxx- VPFT (ours) 75.4 79.5 59.0 64.0 33.8 52.2 56.1 65.2   xxx- VPRL (ours) 91.6 93.2 74.5 77.6 75.8 83.8 80.6 84.9
Models. To explore visual planning without any language influence as confounders and enables a clean investigation, we select models trained exclusively on visual data without any exposure to textual data during pretraining. For visual planning, we use the Large Vision Model (LVM-7B) [4] as the backbone, which is only trained on image sequences and videos. We train the model with 1) supervised fine-tuning over golden planning trajectory (VPFT) and 2) two-stage reinforcement learning (VPRL), resulting in two system variants with visual planning. For RL training, we start with a rule-based parsing function as the dynamics interpreter to parse the image pairs to actions, and a heuristic progress estimator, with details enclosed in Appendix B.3.
For baselines, to facilitate parallel comparison for language-based planning, we adopt Qwen 2.5-VL-Instruct [3], on both inference-only (Direct [^1] Direct denotes answer prediction without being instructed to conduct intermediate reasoning. and CoT) and post-training settings (SFT and RL), trained on the same data as the visual planner. We further evaluate multimodal reasoning performance of proprietary models with Gemini 2.0 Flash [29] and advanced thinking model Gemini 2.5 Pro [14]. Full training details, model versions, and hyperparameters are provided in Appendix B.4.
@@
  • Exact Match (EM) is defined as $\mathrm{EM}=\max_{m\in\{1,\dots,M\}}\prod_{j=1}^{n}\mathbb{I}(\hat{v}_{j}=v^{(m)}_{j})$, evaluating whether $\hat{\mathcal{T}}$ coincides with any $\mathcal{T}^{(m)}\in\mathcal{O}$. EM requires the entire trajectory to be valid and of minimal length, and accepts all optimal solutions rather than a single reference.
  • Progress Rate (PR) is defined as $\mathrm{PR}=\max_{m\in\{1,\dots,M\}}\frac{1}{n}\sum_{j=1}^{n}\left[\prod_{k=1}^{j}\mathbb{I}(\hat{v}_{k}=v^{(m)}_{k})\right$], measuring the ratio of consecutive correct steps (valid forward moves) from the start that align with at least one optimal trajectory. PR thus provides a softer signal than EM, capturing the model’s ability to make meaningful progress towards a full solution.
Figure 3: Illustration of each task with generated visual planning traces from LVM, covering different types of actions (optimal, non-optimal and invalid). More cases can be found in Appendix C.6.
Table 2: Performance of text-based planning variants on FrozenLake. See Table 7 in Appendix C.2 for the full results.
@@
Textual planning falls short in both proprietary models and open-sourced tuning baselines. Table 1 shows that proprietary models yield average EM below 50% and PR only marginally above 50% at best, underscoring the challenges these tasks pose for current models despite being intuitive for humans. On the other hand, while task-specific training provides partial improvement, the overall performance of fine-tuned textual planners remains unsatisfactory, through either directly generating planned actions (SFT in Table 1) or first captioning the image with different textual representations and then generating answers (Table 2). We also observe that, unlike the notable gains of RL in the pure language domain [17], RL yields limited performance gains when applied to text-based planning with multimodal inputs. Table 2 shows that when using progress reward as in VPRL or directly using the Progress Rate metric as the outcome reward, none of the variants surpasses the SFT baseline. We attribute the bottleneck of language-based planning with SFT and RL to the modality gap, which leads to inaccuracies in grounding visual information into text and thereby constrains performance. Further discussion is provided in Section 4.
@@
Visual planning achieves better performance than textual baselines via RL. While supervised fine-tuning (VPFT) achieves performance comparable to text-based SFT, it remains constrained by imitation and limited exposure to diverse trajectories. By contrast, our two-stage reinforcement learning framework (VPRL) substantially improves the planning capability, achieving the strongest overall results. After Stage 2 optimization, the model attains near-perfect accuracy on FrozenLake (91.6% EM, 93.2% PR) and maintains strong performance on more complex Maze and MiniBehavior tasks, outperforming VPFT by over 20% on average. As expected, the improvement is fully driven by outcome-based optimization in Stage 2, as Stage 1 alone yields near-random behavior (Table 10 in Appendix C.6). Unlike VPFT, which mainly fits the training distribution, VPRL enables exploration of diverse actions and learning from their consequences through reward-driven updates, thereby capturing underlying planning rules and achieving stronger performance.
@@
VPRL shows robustness with scaling complexity. The advantage of RL also holds when we study the performance of different methods with respect to task difficulties, where a larger grid usually relates to higher difficulties. In Figure 6, as the grid size increases from $3\times 3$ to $6\times 6$ in the FrozenLake environment, Gemini 2.5 Pro’s EM score drops sharply from 98.0% to 38.8%. In comparison, our visual planners not only maintain higher accuracy at all grid sizes but also exhibit a much flatter performance curve. Similarly, VPRL demonstrates even greater stability than VPFT, with EM remaining at 97.6% on $3\times 3$ grids and still achieving 82.4% on $6\times 6$, indicating strong robustness. We observe similar trends in other tasks; see Appendix C.3 for other tasks.

### v3

@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; Anil et al., 2023) have demonstrated strong capabilities in language understanding and generation, as well as growing competence in complex reasoning, enabled by their chain-of-thought reasoning abilities (Wei et al., 2022b). Building on these advances, recent work extends LLMs to support multiple modalities, yielding so-called Multimodal Large Language Models (MLLMs) (Reid et al., 2024; Hurst et al., 2024): they incorporate visual embedded information at the input to tackle a broader spectrum of tasks, such as visual spatial reasoning (Liu et al., 2023; Li et al., 2024a) and navigation (Gu et al., 2022; Li et al., 2024b). However, despite their multimodal inputs, these methods perform reasoning purely in the text format during inference, from captioning visual content (Hao et al., 2025) to generating verbal rationales (Zhang et al., 2024b).
Building on this observation, we argue that performing multimodal reasoning only in the text pathway may not always offer the most intuitive or effective strategy, particularly for tasks that depend heavily on visual information and/or are ‘vision-first’ by design. Indeed, recent results from multimodal benchmarks (Roberts et al., 2025; Li et al., 2024a; Chen et al., 2024; Cheng et al., 2025) offer growing evidence that purely language-based reasoning falls short in certain domains, particularly those involving spatial, geometric, or physical dynamics (Zhang et al., 2025a). Such reliance on grounding visual information into text before reasoning introduces a modality gap that hinders the model’s ability to capture visual features and state transitions. This highlights a potential shortcoming of current MLLMs: while they process image inputs, they do not naturally “think” in images. For instance, tasks such as planning a route through a maze, designing the layout of a room, or predicting the next state of a mechanical system are often better served by visual representations, as verbal descriptions may be less effective and struggle to accurately capture complex spatial reasoning relationships. These examples suggest a broader question, which we aim to tackle in this work: can models directly plan in non-verbal modalities, such as images, without being mediated by text?
@@
Cognitive science also offers compelling motivation for this question (Moulton & Kosslyn, 2009). Dual Coding Theory (Paivio, 1991) proposes that human cognition operates through both verbal and nonverbal channels, each capable of independent representational and inferential processes. Recent work on MLLMs incorporates interleaved text and images as reasoning steps (Hu et al., 2024; Li et al., 2025). However, they still remain fundamentally text-driven and rely on tool-based visualizations as auxiliary information for reasoning traces, with reasoning still mainly embedded in verbal traces. For instance, Visual Sketchpad (Hu et al., 2024) employs external tools to generate sketches as visual aids, and MVoT (Li et al., 2025) generates per-step visualizations from language-based actions but still reasons in text for decision-making. As such, a truly visual-only reasoning paradigm that avoids any text-based reasoning proxies remains underexplored.
@@
In this work, we propose a new paradigm, Visual Planning, where reasoning is structured as a sequence of images, but without the mediation of language. To the best of our knowledge, this is the first attempt to investigate whether models can achieve planning purely through visual representations. Rather than generating textual rationales and answers, our approach produces step-by-step visualizations that encode planning or inference steps directly in images. As a pioneering exploration, it circumvents the modality mismatch that occurs when visual problems must be forced into explanations in verbal form, reinforces state transitions, and provides a new trackable interface for tasks like navigation (Li et al., 2024a), and visual problem-solving (Hao et al., 2025).
@@
Specifically, we explore this paradigm using the Large Vision Model (LVM) (Bai et al., 2024) trained exclusively on images and video frames with no textual data. This design choice removes potential confounders introduced by language-based supervision and enables a clean investigation of whether models can reason purely within the visual modality. Motivated by the success of reinforcement learning in acquiring reasoning capabilities within the language modality (Guo et al., 2025a) and its strong generalization performance (Chu et al., 2025), we propose Visual Planning via Reinforcement Learning (VPRL), a novel two-stage reinforcement learning framework empowered by GRPO (Shao et al., 2024) for visual planning. It involves a distinct initializing stage for encouraging the exploration of the policy model in the given environment, which is then followed by reinforcement learning with a progress reward function.
@@
We validate the feasibility of our paradigms on grid-based navigation as a representative of spatial planning tasks, including Maze (Ivanitskiy et al., 2023), FrozenLake (Wu et al., 2024b), and MiniBehavior (Jin et al., 2023), where one agent is requested to navigate to a target location successfully without violating environment constraints. Our experiments reveal that the visual planning paradigm substantially surpasses the traditional textual reasoning method by supervised fine-tuning (SFT), achieving 27% higher average exact-match rate. In addition to better performance, our novel method VPRL exhibits stronger generalization to out-of-distribution scenarios than the SFT method in the visual planning paradigm (VPFT). To the best of our knowledge, we are the first to apply RL to image generation in the context of planning, with contributions as follows:
@@ 2.1 The Visual Planning Paradigm
The majority of prior visual reasoning benchmarks (Goyal et al., 2017; Akula et al., 2021; Yue et al., 2024) can be and is typically tackled by grounding the visual information in the textual domain (Gurari et al., 2018; Peng et al., 2024; Zhang et al., 2024a), followed by a few steps of textual reasoning. However, once the visual content is mapped to text (e.g., object names, attributes, or relations), the problem gets reduced to a language reasoning task, where the reasoning is carried out by the language model, even without reflecting any information from the visual modality.
@@ 2.2 Reinforcement Learning for Large Vision Models
Reinforcement learning (RL) has shown notable advantages in improving the generalization of autoregressive models by optimizing with sequence-level rewards beyond token-level supervision signals (Chu et al., 2025). In autoregressive image generation, an image is represented as a sequence of visual tokens. Inspired by the success of RL in language reasoning (Guo et al., 2025a), we introduce an RL-based training framework for visual planning empowered by Group Relative Policy Optimization (GRPO) (Shao et al., 2024). It leverages the transitions between visual states to compute the reward signals while verifying the constraints from the environments. To enforce the policy model that generates valid actions with diverse exploration during the RL process, we then propose a novel two-stage reinforcement learning framework for visual planning. In Stage 1, we first apply supervised learning to initialize the policy model with random trajectories. Model’s visual planning is then optimized by the RL training in Stage 2.
@@
The dynamics interpreter $\mathcal{D}$ evaluates the transitions $a\in\mathcal{A}\cup\mathcal{E}$ for validity, which, by implementation, can be a dynamics model (Qiu et al., 2025) or a rule-based system to elicit actions from state pairs, or a neural model as holistic validator that judges transitions without explicitly inferring actions. The progress estimator $P(v)$ quantifies progress by estimating the remaining steps or effort required to reach the goal from each visual state. By comparing the agent’s current and predicted state, we partition the generated candidate states $\mathcal{A}\cup\mathcal{E}$ into three disjoint subsets:
@@ 3 Experiments and Results
Tasks. To evaluate our proposed visual planning paradigm, we select representative tasks where planning can be expressed and executed entirely in the visual modality. We focus on tasks where state transitions are visually observable, distinguishing them from language-centric tasks like code generation (Lai et al., 2023) or traditional visual question answering. This design allows us to analyze planning behavior without relying on textual rationales or symbolic outputs. To compare visual planning with language-based reasoning, we experiment with 3 visual navigation environments: FrozenLake (Wu et al., 2024b), Maze (Ivanitskiy et al., 2023), and MiniBehavior (Jin et al., 2023). All of them can be solved in both modalities, which enables a direct parallel comparison of pros and cons between visual planning and language reasoning strategies.
  • FrozenLake: It is initially proposed by Wu et al. (2024b) and implemented with Gym (Brockman, 2016). It simulates a grid-based frozen lake, where the agent is supposed to start from the designated position and find its way to the destination safely without falling into the ‘holes’.
@@
We construct synthetic datasets for the tasks with varying levels of complexity in patterns and environments. Details on data collection and implementation are provided in Appendix E.1.
Table 1: Performance of the closed- and open-source models on FrozenLake, Maze, and MiniBehavior. VPRL performs consistently the best (bold) across all tasks. denotes the post-trained model.  represents texts and  represents images. The last column Avg. reports the average performance across three tasks.
Model Input Output FrozenLake Maze MiniBehavior Avg. EM (%) PR (%) EM (%) PR (%) EM (%) PR (%) EM (%) PR (%) Closed-Source Model Gemini 2.0 Flash   xxx- Direct + 21.2 47.6 8.3 31.4 0.7 29.8 10.1 36.3   xxx- CoT + 27.6 52.5 6.9 29.8 4.0 31.2 12.8 37.8 Gemini 2.5 Pro (think) + 72.0 85.0 21.5 35.5 37.6 59.9 43.7 60.1 Open-Source Model Qwen 2.5-VL-Instruct-7B   xxx- Direct + 1.2 15.0 0.6 14.5 0.3 9.8 0.7 13.1   xxx- CoT + 8.2 29.1 2.3 15.2 0.5 14.7 3.7 19.7   xxx- SFT + 68.6 84.4 60.9 70.3 31.3 56.1 53.6 69.9 LVM-7B   xxx- VPFT (ours) 75.4 79.5 59.0 64.0 33.8 52.2 56.1 65.2   xxx- VPRL (ours) 91.6 93.2 74.5 77.6 75.8 83.8 80.6 84.9
Models. To explore visual planning without any language influence as confounders and enables a clean investigation, we select models trained exclusively on visual data without any exposure to textual data during pretraining. For visual planning, we use the Large Vision Model (LVM-7B) (Bai et al., 2024) as the backbone, which is only trained on image sequences and videos. We train the model with 1) supervised fine-tuning over golden planning trajectory (VPFT) and 2) two-stage reinforcement learning (VPRL), resulting in two system variants with visual planning. For RL training, we start with a rule-based parsing function as the dynamics interpreter to parse the image pairs to actions, and a heuristic progress estimator, with details enclosed in Appendix E.3.
For baselines, to facilitate parallel comparison for language-based planning, we adopt Qwen 2.5-VL-Instruct (Bai et al., 2025), on both inference-only (Direct [^1] Direct denotes answer prediction without being instructed to conduct intermediate reasoning. and CoT) and post-training settings (SFT and RL), trained on the same data as the visual planner. We further evaluate multimodal reasoning performance of proprietary models with Gemini 2.0 Flash (Kampf & Brichtova, 2025) and advanced thinking model Gemini 2.5 Pro (Gemini, 2025). Full training details, model versions, and hyperparameters are provided in Appendix E.4.
@@
  • Exact Match (EM) is defined as $\mathrm{EM}=\max_{m\in\{1,\dots,M\}}\prod_{j=1}^{n}\mathbb{I}(\hat{v}_{j}=v^{(m)}_{j})$, evaluating whether $\hat{\mathcal{T}}$ coincides with any $\mathcal{T}^{(m)}\in\mathcal{O}$. EM requires the entire trajectory to be valid and of minimal length, and accepts all optimal solutions rather than a single reference. Here, the equality $\hat{v}_{j}=v^{(m)}_{j}$ refers to whether the two states can be reached from their respective previous states by applying the same action. This means that the comparison is made at the level of environment transitions rather than a pixel-wise match between images. In other words, two states are treated as the same if they represent the same underlying configuration, even when their pixel values are not identical.
  • Progress Rate (PR) is defined as $\mathrm{PR}=\max_{m\in\{1,\dots,M\}}\frac{1}{n}\sum_{j=1}^{n}\left[\prod_{k=1}^{j}\mathbb{I}(\hat{v}_{k}=v^{(m)}_{k})\right$], measuring the ratio of consecutive correct steps (valid forward moves) from the start that align with at least one optimal trajectory. PR thus provides a softer signal than EM, capturing the model’s ability to make meaningful progress towards a full solution. The same state equality is applied as in EM.
Figure 3: Illustration of each task with generated visual planning traces from LVM, covering different types of actions (optimal, non-optimal and invalid). More cases can be found in Appendix F.6.
Table 2: Performance of text-based planning variants on FrozenLake. See Table 7 in Appendix F.2 for the full results.
@@
Textual planning falls short in both proprietary models and open-sourced tuning baselines. Table 1 shows that proprietary models yield average EM below 50% and PR only marginally above 50% at best, underscoring the challenges these tasks pose for current models despite being intuitive for humans. On the other hand, while task-specific training provides partial improvement, the overall performance of fine-tuned textual planners remains unsatisfactory, through either directly generating planned actions (SFT in Table 1) or first captioning the image with different textual representations and then generating answers (Table 2). We also observe that, unlike the notable gains of RL in the pure language domain (Guo et al., 2025a), RL yields limited performance gains when applied to text-based planning with multimodal inputs. Table 2 shows that when using progress reward as in VPRL or directly using the Progress Rate metric as the outcome reward, none of the variants surpasses the SFT baseline. We attribute the bottleneck of language-based planning with SFT and RL to the modality gap, which leads to inaccuracies in grounding visual information into text and thereby constrains performance. Further discussion is provided in Section 4.
@@
Visual planning achieves better performance than textual baselines via RL. While supervised fine-tuning (VPFT) achieves performance comparable to text-based SFT, it remains constrained by imitation and limited exposure to diverse trajectories. By contrast, our two-stage reinforcement learning framework (VPRL) substantially improves the planning capability, achieving the strongest overall results. After Stage 2 optimization, the model attains near-perfect accuracy on FrozenLake (91.6% EM, 93.2% PR) and maintains strong performance on more complex Maze and MiniBehavior tasks, outperforming VPFT by over 20% on average. As expected, the improvement is fully driven by outcome-based optimization in Stage 2, as Stage 1 alone yields near-random behavior (Table 10 in Appendix F.6). Unlike VPFT, which mainly fits the training distribution, VPRL enables exploration of diverse actions and learning from their consequences through reward-driven updates, thereby capturing underlying planning rules and achieving stronger performance.
@@
VPRL shows robustness with scaling complexity. The advantage of RL also holds when we study the performance of different methods with respect to task difficulties, where a larger grid usually relates to higher difficulties. In Figure 6, as the grid size increases from $3\times 3$ to $6\times 6$ in the FrozenLake environment, Gemini 2.5 Pro’s EM score drops sharply from 98.0% to 38.8%. In comparison, our visual planners not only maintain higher accuracy at all grid sizes but also exhibit a much flatter performance curve. Similarly, VPRL demonstrates even greater stability than VPFT, with EM remaining at 97.6% on $3\times 3$ grids and still achieving 82.4% on $6\times 6$, indicating strong robustness. We observe similar trends in other tasks; see Appendix F.3 for other tasks.